OpenStudy (anonymous):

Find -3a^2 - b^3 + 3c^2 - 2b^3 if a = 2, b = -1, and c = 3 ^=exponent Please help I'm way beyond confused

OpenStudy (midhun.madhu1987):

The same method as u have done bfore.. just plug in the value of a,b and c. make a try.. i'll help if there are any issues....

OpenStudy (anonymous):

ok -3(2)^2-(-1)^3+3(3)^2-2(-1)^3=

OpenStudy (anonymous):

Thats simplfyed soooo

OpenStudy (anonymous):

-3x2= -6^2= -36 -1^3=-1 -36-(-1)= -37 3x3=9^2= 81 -37+81=44 -1= 43 :) is that correct

OpenStudy (phi):

You have to be more careful -3(2)^2-(-1)^3+3(3)^2-2(-1)^3= -3(2)^2 means -3*2*2 can you simplify that ?

OpenStudy (anonymous):

36positive or 12?

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (phi):

do it in steps -3*2*2 first do -3*2 what do you get ?

OpenStudy (anonymous):

-6

OpenStudy (phi):

now do -6 * 2 what do you get ?

OpenStudy (anonymous):

-12?

OpenStudy (phi):

yes -3*2*2= -12 do you see how to do it ? could you do it yourself ?

OpenStudy (anonymous):

can you still be here just in case i do something wrong

OpenStudy (phi):

ok

OpenStudy (phi):

can you try -(-1)^3 first, it means - (-1)*(-1)*(-1) -1*-1 is what ? (remember 2 minus makes a plus)

OpenStudy (anonymous):

-3(2)^2-(-1)^3+3(3)^2-2(-1)^3 -12-3+81-1=

OpenStudy (anonymous):

is that correct??

OpenStudy (anonymous):

so far

OpenStudy (phi):

we should do each term, one at a time, until you get expert. -(-1)^3 first, it means - (-1)*(-1)*(-1) -1*-1 is what ?

OpenStudy (anonymous):

ohh -2 :)

OpenStudy (anonymous):

and -2*-1= -3

OpenStudy (phi):

-1 * -1 is negative time negative. remember: 2 minus make a plus try again.

OpenStudy (anonymous):

2 then not -2

OpenStudy (phi):

when you MULTIPLY, and you see two minus signs, cross them off change -1 * -1 to 1 * 1 (because we are multiplying, and when we see two minus signs, we can erase them... try to remember that rule)

OpenStudy (phi):

when you see -1 * -1, make it 1*1 what is 1 times 1 ?

OpenStudy (anonymous):

1

OpenStudy (phi):

- (-1)*(-1)*(-1) we do the -1 * -1 and it is 1 - ( 1* -1) 1*-1 is -1 -(-1) what is that ?

OpenStudy (anonymous):

huhhh :((( what am I subtracting it from

OpenStudy (phi):

we are doing the 2nd term in -3(2)^2 -(-1)^3 +3(3)^2-2(-1)^3 which is -(-1)^3 if that seems confusing, change the - to -1* -1*(-1)^3 now write (-1)^3 as -1*-1*-1 and the problem is -1*-1*-1*-1 if you write it this way (-1 *-1) * (-1 * -1) can you simplify that ?

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

i'm starting to think its just 1 but thats not what it is?

OpenStudy (phi):

(-1 *-1) is 1 (remember the rule minus times minus, make it positive 1*1 and 1 times 1 is 1 that means (-1 *-1) * (-1 * -1) becomes 1 * 1 and 1 times 1 is 1 so the answer is -(-1)^3 = 1

OpenStudy (anonymous):

thats it is that the whole answer to the whole ? or just that part?

OpenStudy (phi):

just the second term so far we have -12 + 1 now the 3rd term 3(3)^2 first, rewrite the (3)^2 using multiply. can you do that ?

OpenStudy (anonymous):

9 righttt? and then overall its 81

OpenStudy (phi):

3(3)^2 the (3)^2 means 3*3 3(3)^2 means 3* (3)^2 or 3*3*3 that is not 81

OpenStudy (anonymous):

27 :) how do i not remember that

OpenStudy (phi):

ok now we have -12 + 1 + 27 now the last term -2(-1)^3 remember this means -2 * (-1)^3 and (-1)^3 means what ?

OpenStudy (anonymous):

3?

OpenStudy (phi):

exponents mean multiply by itself (*not* add)

OpenStudy (anonymous):

ya -1*-1*-1=3 right because if there are negatives its a positive

OpenStudy (phi):

you added. multiply is different. and the rule is 2 negatives make a positive. that means you can erase *two* of the minus signs, but that leaves one still there. so first, use the 2 minus make a plus to change -1*-1*-1 to 1*1*-1 now multiply (not add)

OpenStudy (anonymous):

0???

OpenStudy (anonymous):

or -1

OpenStudy (phi):

1*1 is ?

OpenStudy (anonymous):

1

OpenStudy (phi):

and 1*-1 is ?

OpenStudy (phi):

-1*-1*-1 is (-1 * -1) * -1 and that is 1 * -1 and that is -1

OpenStudy (anonymous):

so its -1 ok :) that makes since then

OpenStudy (phi):

now we can do -2 * (-1)^3 which is -2 * -1 what is that ?

OpenStudy (anonymous):

2 :)

OpenStudy (phi):

so now you have -12 + 1 + 27 + 2

OpenStudy (anonymous):

18 :)

OpenStudy (anonymous):

@midhun.madhu1987